Help with thermodynamics basics? Due at midnight, please help!

refer to a cup which contains water at room temper
Answer #1

16.
Since water doesn't boil, temperature will never be constant.
Also initially, temperature will initially rise faster and later will become slow as it is reaching equilibrium
Answer: E

17.
Q = m*C* delta T
Is Q is doubled, delta T will double
Answer: B

18.
Q = m*C* delta T
Is m is halfed, delta T will double
Answer: B

19.
Q = m*C* delta T
Is C is halfed, delta T will double
Answer: B

Only 4 subparts at a time please